Simplify 52(x-1)+5x+2=0.04

25

25


1
Expert's answer
2021-02-24T07:56:03-0500

"5^{2(x-1)}+5^{x+2}=0.04\newline
5^{-2}5^{2x}+5^2*5^{x}=5^{-2}"

"5^{2x}+5^4*5^{x}-1=0"

"y =5^x;y>0"

"y^2+625y-1=0"

"y= -312.5+\sqrt{97657.25}"

"x=\log_5(-312.5+\sqrt{97657.25})"

Answer:"x=\log_5(-312.5+\sqrt{97657.25})"